# DiLi: A Lock-Free Asynchronously Distributable Linked List

### v1

@@ 6.2 Conditional lock-freedom of Client Operations
Conditional lock-freedom depends on the fact that any client operation utilizes atmost four threads (three when there is no maintenance) and that the number of instructions executed by a delegating thread is bounded (proved in Section˜D.2). When there is no background operation, being similar to a Harris List, a successful Remove causes at most two updates to a sublist – one for marking (by CAS in Line 2) and one for delinking (by CAS in Line 2). A successful insert causes at most one update (by CAS in Line 3). The number of retries in Search is limited by the number of marked nodes to be delinked from concurrent Removes. Hence, no operation requires more retries than the number of concurrent updates [14].
@@
During a Split operation, the search traversal (as denoted by the red lines in Algorithm˜2) is only enhanced by an earlier truncation from the introduced subtail. An earlier truncation that occurs exactly at the item where a concurrent insert is taking place shortcuts to the next subhead (red lines of Algorithm˜3, as the insertion is now expected to happen on the subhead instead of the item where the split is taking place, in order to preserve the key range property of the split sublists. By having zero additional search traversals for find and remove, and exactly one additional search traversal, the Harris List based operations that were lock-free when there is no split operation, are also lock-free during a Split operation.
@@
During a Switch of a sublist, until the sublist entries of all servers are updated with the latest subhead, there will be an additional delegation to the right machine using the blue lines of Algorithm˜2 and Algorithm˜3 where the stCt values are checked to be negative. Thus, any concurrent client operation on the sublist during this time suffers from exactly one additional network hop and search traversal on the sublist. Since the search traversal and client operations were lock-free in the absence of a Switch, the client operations are lock-free even during the Switch.
7 Empirical Evaluation
7.1 Implementation
@@
We implemented DiLi in C++11 in the Ubuntu 24.04.1 LTS operating system and utilized Google’s RPC framework (gRPC) v1.66 [13]. The RPC requests are processed asynchronously based on thread availability, while the requester of any RPC synchronously waits for a response. The only situation a requester instead requests asynchronously is to send replication of updates whose responses are processed as asynchronous callbacks by a separate group of threads.
@@
Each machine that serves DiLi is assigned an initial key range to serve the list, chosen naively by a range partitioning on the key range of the list. We also implement a naive load balancer that utilizes the Split, Move, and Switch algorithms. The load balancer is a separate thread spawned in each machine to repeatedly traverse through all sublists held by the machine and to find sublists that are bigger than a threshold of 125 in size, and to use Split roughly in the middle to create smaller sublists for quicker traversal. Additionally, a rough size of the list is measured by taking the individual sum of the sublist sizes and exchanged between machines to make a load balancing decision. A decision to move is made when a machine holds more than 110% of its assigned load, and invokes Move on one of its sublists to a machine with the least load. In this way, the Split invocations control the amount of linear traversal on the sublist for client operations, and the Move invocations roughly balance the number of items present in each machine. Using the same gRPC framework, we also implemented the standard lock-free skip list [11] in C++, and the Harris Linked list for comparison. All implementations can be found in this anonymized repository [23].
7.2 Setup and Methodology
@@
We perform two main experiments to evaluate the performance of DiLi: 1. A single-machine comparison of DiLi with a lock-free skip list and the Harris Linked list. 2. A Distributed Scalability test to show that DiLi is scalable as we add more number of machines.
@@
Both experiments utilize Yahoo! Cloud Serving Benchmark (YCSB) workloads [8] to simulate workloads of a Zipfian distribution. We do a load of 1M keys into the list and then perform the experiment for a workload of 2M operations, for three different workload read-write proportions : 10% read, 50% read and 90% read. The write proportion is equally split between inserts and removes, to keep the size of the list roughly the same.
@@
The experiments were performed using C7i instances from Amazon Web Services (AWS) as the server machines for the list. The instance features an Intel(R) Xeon(R) Platinum 8488C (2.4GHz, 105MB L3 cache), with each core supporting 2 way hyperthreading. The single-machine comparison utilized an 8 core instance of it with 32 GB of RAM, and the distributed list comparison had multiple 2 core instances, each with 8 GB of RAM.
@@
The results include the throughput measured as the number of operations completed per second. To obtain the maximum throughput achieved by these algorithms, we perform several executions of the same experiment, varying the number of clients through a separate large compute instance, until the server machines reach close to the maximum CPU utilization ($\sim 100\%$ as monitored on AWS CloudWatch). A dry run of 2 iterations is first performed before running either experiment to warm up the servers. Then each experiment is repeated 3 times and the average of those executions is used to compute the reported results.
7.3 Results and Discussion
@@
(a) Single Machine Performance of DiLi as compared to skip lists and Harris List.
@@
(b) Distributed Scalability of DiLi. Here DiLi_x stands for DiLi with x number of similar servers.
Single-Machine Comparisons: The results are in Figure˜3(a). As expected, the use of Split in the background to generate smaller sublists permits significantly faster search compared to Harris List. In fact, DiLi provides a comparable performance to skip lists, that are designed to permit faster traversals on a list. A key parameter for skip lists is the maximum number of levels that it permits. For 1M items, 25 levels are sufficient to have a $\log(n)$ expected traversal time. Hence we test it for multiple levels (10, 25 and 100) around it. From this result, we observe that the ability to horizontally scale across multiple machines has not created a loss in throughput if DiLi is run on a single machine. DiLi provides additional benefits over a skip list. Specifically, the memory usage after the 1M item load was around 370 MB for the 25-level skip list, whereas in DiLi it was only 170 MB. This is expected because the memory usage of a skip list grows by an additional factor of the number of levels in the skip list. We do not provide client operation latency results here, as we observed that the majority of the latency for skip lists and DiLi were caused only by the round-trip delay between the client and the server.
Distributed Scalability: We evaluate the scalability of DiLi by comparing the maximum throughput with 2, 4, 6 and 8 servers (cf. Figure˜3(b)) for each of the aforementioned workloads. Interestingly, we not only observe our claim that DiLi scales linearly for every workload, but we also observe that the scaling differs based on the type of workload. This illustrates how our naive implementation of the load balancer affects the rate of growth in performance. This calls for future work on how the Split and Move interface could be properly utilized to optimize performance for a chosen workload.
@@
Performance of Background Operations: We additionally performed a minor experiment in a 2 machine setup, where the first machine serves the entire key range for the list, and a second machine starts without owning any part of the key range. We performed the 1M key load onto this setup, and set the load balancer to transfer sublists to the idle server dynamically during an insert load, until both machines have comparable number of items. During this test, we measured the latency of the Split and Move operations (as depicted in Appendix˜C). The average latency to Split was 2.65 ms and the average latency to Move a sublist was 51.53 ms. Thus, we empirically verify that the asynchronous operations can terminate quickly even during an insert dominated workload.
8 Related Work
@@
Linked lists are fundamental data structures. A sequential linked list is easy to implement. However, a concurrent link list is challenging to design, and hence has been studied extensively for decades. Consequently, there exist many lock-free implementations – [30, 14, 9]; and a few wait-free implementations – [31, 29]. The first lock-free implementation using the atomic CAS primitive (for a non-distributed list) is due to Valois [30]. Harris [14] gives a lock-free ordered list implementation. Among lock-free lists with no indexing and versioning, Harris List has been the state-of-the-art. Heller et al. [16] give a lock-based linked list design with a wait-free lookup operation. Our linear traversal of client operations resembles that of a Harris list. A wait-free ordered linked list implementation is given by Timnat et al. [29].
@@
Some papers focused on optimizing the cost of operations in addition to correctness and progress guarantees established in the aforementioned papers. An example is a lock-free list implementation due to Fomitchev and Ruppert [9], with worst-case linear amortized cost. [14] and [9] both utilize a $\tt{HelpMarked}$ in their read operations. They perform de-linking of marked nodes during any traversal. Zhang et al.[31] provided an unordered linked list where insertions always occur in the head position of the linked list.
@@
A concurrent lock-free doubly-linked list implementation is given by Attiya and Hillel [4], using a double CAS operation. This doubly-linked list does not have a read operation. The same is implemented using only CAS by Sundell and Tsigas [27].
@@
The aforementioned implementations typically consider the sequences of entries in the linked list to reside in different chunks of memory. To take advantage of entries that are instead on the same chunk, Braginsky and Petrank [6] developed a concurrent linked list that maintains in each memory chunk, a certain number of entries within a defined minimum and maximum. Their approach shares some similarities with our key range component of a sublist. Such chunk mechanisms have also been explored in key-value stores that use a versioned linked list based search structure [5]. We note that this optimization is also applicable to the sublists of DiLi. We do not add this to our implementation to have a fair comparison with skip lists, which cannot have such an optimization across its levels.
@@
The versioned linked list is a classic technique that arose from the need for read-only transactions in databases [20]. Since we only support basic operations insert, remove, and find of a key in this paper, we lack the need for multi-versioning and hence do not compare our performance with versioned lists. Such databases [5, 21] also use lock-free skip lists to boost index search performance. These lock-free skip lists ([28, 11, 10]), can be replaced with DiLi, as it offers comparable performance and horizontal scalability via distribution.
@@
Another well-celebrated direction is to use a universal construction approach, for example, Herlihy [17], to convert sequential implementations to wait-free concurrent implementations. However, universal construction adds overhead compared to direct implementations.
@@
The aforementioned literature are all implementations on a single machine. Although there exist lock-based implementations of a distributed B-Tree in the literature [26], with regard to nonblocking distributed implementations on multiple machines, the literature is rare. The most related work is due to Abe and Yoshiday [2] where a doubly linked distributed list is constructed. The presented strategy uses conflict detection and sequence numbers (with some assumptions) to break symmetries. It guarantees atomic insert and delete operations and non-blocking lookup operations, but the implementation is only obstruction-free.
@@
9 Concluding Remarks
In this paper, we focussed on developing DiLi, a distributed linked list data structure. The goal of this data structure is to preserve the client properties of Harris List (specifically, the lock-free nature) while providing horizontal scaling. Specifically, as the size of the data structure grows and as the number of operations performed on it in a given time increases, the data structure needs to be scaled to handle it.

### v2

@@ 6.2. Conditional lock-freedom of Client Operations
Conditional lock-freedom depends on the fact that any client operation utilizes atmost four threads (three when there is no maintenance) and that the number of instructions executed by a delegating thread is bounded (proved in Appendix (Me, 2025b)). When there is no background operation, being similar to a Harris List, a successful Remove causes at most two updates to a sublist – one for marking (by CAS in Line 1) and one for delinking (by CAS in Line 1). A successful insert causes at most one update (by CAS in Line 1). The number of retries in Search is limited by the number of marked nodes to be delinked from concurrent Removes. Hence, no operation requires more retries than the number of concurrent updates (Harris, 2001).
@@
During a Split operation, the search traversal (as denoted by the red lines in Algorithm 1) is only enhanced by an earlier truncation from the introduced subtail. An earlier truncation that occurs exactly at the item where a concurrent insert is taking place shortcuts to the next subhead (red lines of Algorithm 1, as the insertion is now expected to happen on the subhead instead of the item where the split is taking place, in order to preserve the key range property of the split sublists. By having zero additional search traversals for find and remove, and exactly one additional search traversal, the Harris List based operations that were lock-free when there is no split operation, are also lock-free during a Split operation.
@@
During a Switch of a sublist, until the sublist entries of all servers are updated with the latest subhead, there will be an additional delegation to the right machine using the blue lines of Algorithm 1 and Algorithm 1 where the stCt values are checked to be negative. Thus, any concurrent client operation on the sublist during this time suffers from exactly one additional network hop and search traversal on the sublist. Since the search traversal and client operations were lock-free in the absence of a Switch, the client operations are lock-free even during the Switch.
7. Empirical Evaluation
7.1. Implementation
@@
We implemented DiLi in C++20 in the Ubuntu 24.04.1 LTS operating system and utilized Google’s RPC framework (gRPC) v1.72 (Google, 2016). The RPC requests are processed asynchronously based on thread availability, while the requester of any RPC synchronously waits for a response. The only situation a requester instead requests asynchronously is to send replication of updates whose responses are processed as asynchronous callbacks by a separate group of threads.
@@
Each machine that serves DiLi is assigned an initial key range to serve the list, chosen naively by a range partitioning on the key range of the list. We also implement a naive load balancer that utilizes the Split, Move, and Switch algorithms. The load balancer is a separate thread spawned that competes with threads performing the list operations in each machine. It repeatedly traverses through all sublists held by the machine to find sublists that are larger than a threshold of size (various values were experimented), and to use Split roughly in the middle to create smaller sublists for quicker traversal. Additionally, a rough size of the list is measured by taking the individual sum of the sublist sizes and exchanged between machines to make a load balancing decision. A decision to move is made when a machine holds more than 110% of its assigned load, and invokes Move on one of its sublists to a machine with the least load. In this way, the Split invocations control the amount of linear traversal on the sublist for client operations, and the Move invocations roughly balance the number of items present in each machine. Using the same gRPC framework, we also implemented the Java standard lock-free skip list (Fraser, 2004) in C++20, and the Harris Linked list for comparison. All implementations can be found in this anonymized repository (Me, 2025a).
We have also implemented a variation, where the linked list is packaged as a ‘library’ that can be operated concurrently by multiple threads. We use this variation of skip list and DiLi to evaluate the performance when there is no network communication overhead. The purpose of this experiment is to illustrate that while DiLi is designed to distribute a list across multiple machines, it provides a performance comparable to a skip list when restricted to a single machine.
7.2. Setup and Methodology
@@
We perform three main experiments to evaluate the performance of DiLi: (1) Multi-thread scalability of DiLi when compared to a lock-free skip list. (2) A single-server multi-client comparison of DiLi with a lock-free skip list and the Harris Linked list. (3) A Distributed Scalability test to show that DiLi is scalable as we add more number of machines.
@@
All three experiments modify Yahoo! Cloud Serving Benchmark (YCSB) workload template A (Cooper et al., 2010) to generate workloads of a Zipfian distribution. We load 1M keys into the list and then perform the experiment for a workload of 2M operations for five different workload write proportions : 10%, 25%, 50%, 75% and 90% write. The write proportion is equally split between inserts and removes, to keep the size of the list roughly the same. For the library experiment, we additionally evaluate five different loaded key sizes : 100K, 1M, 2M, 5M and 10M, while performing 2M operations of 50% write workload.
@@
The experiments were performed using C7i instances from Amazon Web Services (AWS) as server machines for the list. The instance features an Intel(R) Xeon(R) Platinum 8488C (2.4GHz, 105MB L3 cache), with each core supporting 2 way hyperthreading and a $\sim 12$ Gigabit network. The library and single-server comparison utilized an 8 core instance of it with 32 GB of RAM, and the distributed list evaluation had multiple 2 core instances, each with 8 GB of RAM.
@@
For the distributed setup, the results include the throughput measured as the number of operations completed per second. To obtain the maximum throughput achieved by these algorithms, we perform several executions of the same experiment, varying the number of clients through a separate large compute instance, until a peak throughput is reached. Note here that the performance is evaluated on a low-latency network. Increased latency would give us different set of results (much less throughput) due to the nature of the RPC framework and TCP connections. A dry run of 2 iterations is first performed before running either experiment to warm up the servers. Then each experiment is repeated 3 times and the average of those executions is used to compute the reported results. In the library setup there is no network overhead that can overshadow the operation latencies. Hence we also report the average operation latency (CPU time) of each operation in this experiment.
(a) Maximum average throughput achieved when the write proportion is varied.
(b) Minimum average latency of operations achieved when the write proportion is varied.
(c) Scaling of skip list and DiLi for a read heavy workload.
(d) Scaling of skip list and DiLi for a write heavy workload.
(e) Maximum average throughput achieved when the key size is varied.
(f) Minimum average latency of operations achieved when the key size is varied.
(g) Scaling of skip list and DiLi for 5 million keys.
@@
(h) Single Machine Performance of DiLi as compared to skip lists and Harris List.
@@
(i) Distributed Scalability of DiLi for 50% Write Workload. Here DiLi_x stands for DiLi with x number of similar servers.
Figure 3. Performance of DiLi in library mode and distributed client-server architecture.
7.3. Results and Discussion
Library Experiments: To evaluate DiLi as a multi-threaded library, we measure its throughput by increasing the number of concurrent threads operating on it and compare it to Fraser’s skip list. We perform this evaluation for (1) varying write proportions; (2) varying key sizes; of the workloads.
Write Proportion: As we set the number of keys to 1 million, we use 20 as the optimal maximum level for skip lists. In Figure 3(a) and Figure 3(b), we observe that the performance of DiLi improves as we reduce the threshold size of a sublist. When the threshold is set to 60, it even outperforms the skip list in read heavy workloads by 5.7% and write heavy workloads by 26.7%. As the write proportion increases, the performance of the skip list drops by about 34%, while DiLi_\_60Sub drops only by 21%. This shows that DiLi is also more stable than a skip list when write proportions are varied. Interestingly, as shown in Figure 3(c) and Figure 3(d), the thread wise scaling of DiLi does not stop at 16 threads, which is the maximum number of concurrent threads supported by the hardware. This is important to note because the additional background thread for executing $\tt Split$} consumes CPU cycles in varied amounts. DiLi outperforms the skip list, even when the background thread causes it to be oversubscribed for the same hardware.
Key Size: We also measure the library performance by varying the key size of the linked list, while keeping the write proportion at 50%. In Figure 3(e), we observe that the performance of DiLi is only comparable to skip lists in high key sizes. DiLi_\_40Sub only marginally outperforms skip list at 5 million items by 3.8% and DiLi_\_60Sub outperforms skip list at 2 million items by 8.847%. We also note that there is no ”one size fits all” for DiLi. DiLi_\_60Sub is 8% has better throughput than DiLi_\_40Sub at 2 million, while being 13% worse at 5 million key size.
Distributed Architecture Experiments: Unlike a multi-threaded library, a distributed client-server model has the bottleneck of the network architecture. We compare single-server DiLi with its Ref overhead to a single-server skip list. Then we evaluated the distributed scalability of DiLi, which is our main contribution.
Single-Server Setup: In a client-server setup, we evaluate skip_20 and DiLi in the presence of a network overhead (using low latency gRPC(Google, 2016) channels). We benchmarked the maximum throughput possible to be 979K using a ping RPC. As a result, this reduces the throughput of any implementation to about one tenth of that of the library experiment. Unlike skip_20, for DiLi, we even add the distribution overhead such as Ref pointer, checks for rerouting a request, keeping a port open to listen to other servers, etc. Interestingly, skip list and DiLi have comparable performance ($\sim$ 7% difference in throughput) across workloads of all write proportions.
Distributed Scalability: The true benefit of DiLi is its ability to be scaled not only by multiple concurrent threads on a powerful machine, but by multiple weaker machines that serve the list. Note here that there is also a performance cost of having a network ‘hop’ (delegation) when a client requests a machine that does not have a key. Since the performance from Figure 3(h) has been similar across all workloads, here we evaluate the list for 50% Write workload, with a key size of 1 million. We set the sublist size as 60, as observed to be optimal from the library experiments, and focus on varying the number of machines. In Figure 3(i), we observe that the performance of DiLi scales linearly with the number of machines.
@@ 7.3. Results and Discussion
Performance of Background Operations: We additionally performed an experiment in a 2 machine setup, where the first machine serves the entire key range for the list, and a second machine starts without owning any part of the key range. We performed the 1M key load onto this setup and set the load balancer to transfer sublists to the idle server dynamically during an insert load, until both machines have comparable number of items. During this test, we measured the latency of the Split and Move operations (as depicted in Appendix(Me, 2025b)). The average latency to Split was 2.65 ms and the average latency to Move a sublist was 51.53 ms. Thus, we empirically verify that the asynchronous operations can terminate quickly even during an insert dominated workload.
7.4. Analysis from the Empirical Evaluation
We observe that DiLi can be used a multi-threaded library that performs better than a skip list in write heavy workloads of key sizes as large as a million. DiLi is able to achieve this by having the overhead of a background thread on the same hardware setup. The sublist threshold size parameter can be adjusted based on the expected key size of the list, to further improve the performance. Moreover, the resulting performance is maintained by only using one background thread for the setup. As we scale the setup to multiple machines, we can also choose to have multiple background threads per machine, and have them look at non-overlapping subsets of the registry for faster Splits as the key sizes get much larger.
In a distributed setup, we observe that the network overhead throttles performance of skip lists. One can reduce this overhead using RDMA networks. Regardless, because of dynamic load balancing, we are able to scale such performance with multiple machines. Note that we only used a simple load balancing logifc implemented from our primitives. Tailoring a load balancing algorithm for various workloads can be a research on its own.
In summary, DiLi provides comparable performance to skip lists when restricted to a single machine, where clients may be on the same machine or communicating via a network. At the same time, DiLi provides horizontal scalability by being able to provide increased performance by adding more servers to maintain the list.
8. Related Work
@@
Linked lists are fundamental data structures. A sequential linked list is easy to implement. However, a concurrent link list is challenging to design, and hence has been studied extensively for decades. Consequently, there exist many lock-free implementations – (Valois, 1995; Harris, 2001; Fomitchev and Ruppert, 2004b); and a few wait-free implementations – (Zhang et al., 2013; Timnat et al., 2012). The first lock-free implementation using the atomic CAS primitive (for a non-distributed list) is due to Valois (Valois, 1995). Harris (Harris, 2001) gives a lock-free ordered list implementation. Among lock-free lists with no indexing and versioning, Harris List has been the state-of-the-art. Heller et al. (Heller et al., 2006) give a lock-based linked list design with a wait-free lookup operation. Our linear traversal of client operations resembles that of a Harris list. A wait-free ordered linked list implementation is given by Timnat et al. (Timnat et al., 2012).
@@
Some papers focused on optimizing the cost of operations in addition to correctness and progress guarantees established in the aforementioned papers. An example is a lock-free list implementation due to Fomitchev and Ruppert (Fomitchev and Ruppert, 2004b), with worst-case linear amortized cost. (Harris, 2001) and (Fomitchev and Ruppert, 2004b) both utilize a $\tt{HelpMarked}$ in their read operations. They perform de-linking of marked nodes during any traversal. Zhang et al.(Zhang et al., 2013) provided an unordered linked list where insertions always occur in the head position of the linked list.
@@
A concurrent lock-free doubly-linked list implementation is given by Attiya and Hillel (Attiya and Hillel, 2006), using a double CAS operation. This doubly-linked list does not have a read operation. The same is implemented using only CAS by Sundell and Tsigas (Sundell and Tsigas, 2008).
@@
The aforementioned implementations typically consider the sequences of entries in the linked list to reside in different chunks of memory. To take advantage of entries that are instead on the same chunk, Braginsky and Petrank (Braginsky and Petrank, 2011) developed a concurrent linked list that maintains in each memory chunk, a certain number of entries within a defined minimum and maximum. Their approach shares some similarities with our key range component of a sublist. Such chunk mechanisms have also been explored in key-value stores that use a versioned linked list based search structure (Basin et al., 2020). We note that this optimization is also applicable to the sublists of DiLi. We do not add this to our implementation to have a fair comparison with skip lists, which cannot have such an optimization across its levels.
@@
The versioned linked list is a classic technique that arose from the need for read-only transactions in databases (Keidar and Perelman, 2015). Since we only support basic operations insert, remove, and find of a key in this paper, we lack the need for multi-versioning and hence do not compare our performance with versioned lists(Basin et al., 2020; Kobus et al., 2022). Databases such as SingleStore(Prout et al., 2022) also use lock-free skip lists to boost index search performance. These lock-free skip lists ((Sundell and Tsigas, 2004; Fraser, 2004; Fomitchev and Ruppert, 2004a)), can be replaced with DiLi, as it offers comparable performance and multi-machine scalability.
@@
Another well-celebrated direction is to use a universal construction approach, for example, Herlihy (Herlihy, 1990), to convert sequential implementations to wait-free concurrent implementations. However, universal construction adds overhead compared to direct implementations.
@@
The aforementioned literature are all implementations on a single machine. Although there exist lock-based implementations of a distributed B-Tree in the literature (Sowell et al., 2012), with regard to nonblocking distributed implementations on multiple machines, the literature is rare. The most related work is due to Abe and Yoshiday (Abe and Yoshiday, 2015) where a doubly linked distributed list is constructed. The presented strategy uses conflict detection and sequence numbers (with some assumptions) to break symmetries. It guarantees atomic insert and delete operations and non-blocking lookup operations, but the implementation is only obstruction-free. There also exists distributed data structures that unlike DiLi, are built for specialized hardware / operating systems. . Alam et al. (Alam et al., 2014) developed a distributed skip list for systems that support FG-MPI(Kamal and Wagner, 2010).The skip list here is built for ranged queries, performs only one operation at a time and has loose consistency guarantees for performance trade-offs using skip-list shortcuts. Cell(Mitchell et al., 2016) is a lock-based distributed B Tree that requires RDMA networks for communication.
@@
9. Concluding Remarks
In this paper, we developed DiLi, a distributed linked list data structure. The goal of this data structure is to preserve the client properties of Harris List (specifically, the lock-free nature) while providing horizontal scaling. Specifically, as the size of the data structure grows and as the number of operations performed on it in a given time increases, the data structure needs to be scaled to handle it.